# Do You Get the Hint? Benchmarking LLMs on the Board Game Concept

### v1

@@ 3.2 Prompting
We implement two prompts for the main experiments (in English). Specifically, we create a static and a dynamic prompt. The static prompt consists of the final set of clues given by Player 1 (similar to a snapshot of the clues as they were at the end of the round), and any incorrect guesses made by the LLM are added. We allow 10 guesses per round. This informs us on how well LLMs can navigate the conceptual space, created by humans, and how well they can combine symbolic clues to infer the correct answer. However, as indicated by prior research, including the interaction between players gives more insights into Theory of Mind capabilities. Specifically, in Concept, Player 2 should keep in mind the strategic intent of Player 1’s choice of clues, and why Player 1 might choose to update their clues given Player 2’s incorrect answers. Therefore, in the dynamic prompt, we prompt iteratively, giving the clues from Player 1 per step (one step = the clues given until the original (human) Player 2 interrupted by making a guess). Then, after each step, we add the incorrect guesses from the human players as well as the LLM’s, and update the clues. In this setting, we allow as many guesses as there were made by humans in the original game. We include the prompts in Appendix B.
@@
For the multilingual prompting, we translated the English prompt to the target language [^6] Each translation was proofread by at least a CEFR B2 (ACTFL Advanced Mid) speaker.. We compared translating the entire prompt to the target language, and keeping the instructions in English on a validation set (Vatsal et al., 2025). Given that the latter yielded better results, on the test set we prompt with the entire prompt in the target language. We include the prompts in Appendix B.
@@ 3.3 Evaluation
We evaluate model performance using precision at k (prec@k), where a prediction is considered correct only if the guessed concept exactly matches the target concept. To ensure fair comparison, we apply extensive normalization to the LLM outputs, including correcting misrepresented characters (e.g., ASCII artifacts in non-English concepts), and removing formatting artifacts such as the prefix ‘**Concept**:‘ when present. For the static prompt setting, we fix k at 10. For the dynamic prompt setting, k is set to the number of guesses made by humans in the corresponding round (see Figure 2).
@@ 4 Results
It is clear that Concept is an easy task for humans. To some extent, for persons who play the game very frequently, there might be a learning effect, but since there are 990 unique concepts per language in the game and our dataset includes 100 unique game plays, we can assume this effect is negligible. Considering all rounds, humans achieve a 92% found rate, and 93% on the filtered rounds. 63% of the rounds were solved within 10 guesses, although the correct answer was often still reached when more guesses were required. As expected, there was a negative correlation (Pearson’s r=0.3r=-0.3) between the number of clues provided and the likelihood of success, which could partially be explained because more difficult concepts requiring more clues. To examine whether performance varied across types of concepts, we categorized rounds by their first (and most important) clue (e.g., wildlife/animal, object/thing/package, building/construction/city, music/song/note), focusing on the top 20 most frequent categories. A Bayesian mixed-effects logistic regression model (GLMM) [^7] https://bambinos.github.io/bambi/notebooks/getting_started.html#generalized-linear-mixed-models was then fit with ‘found’ (i.e., whether the concept was found) as the outcome and category as a predictor [^8] Posterior estimates were obtained with NUTS sampling (4 chains; 1,000 warmup and 1,000 post-warmup draws), and category effects were considered meaningful when their 95% highest-density intervals excluded 0.. The analysis revealed no meaningful differences across categories, indicating that human performance was consistently high regardless of concept type.
In comparison, all LLMs perform significantly worse than humans across the board (see Table 2). Interestingly, in rounds where both humans and the LLM found the correct concept, the LLM overall needs fewer guesses to do so, suggesting greater efficiency once they are on the right track (see Figure 3).
Figure 3: Comparison of how many guesses were needed to arrive at the correct concept, for concepts that are found both by humans and the LLM.
However, in rounds where the correct answer is not found easily, LLMs often get stuck pursuing wrong directions, whereas humans are better at revising their earlier interpretations and still arriving at the correct answer [^9] For models, we observe a strong negative correlation between found rate and number of guesses (Pearson’s $r$ between -0.68 and -0.74), whereas for humans the correlation is weak (r=0.14r=-0.14).. This finding echoes Chollet’s 2019 view of exploration as a key component of intelligence: humans can flexibly revise their hypotheses, but LLMs tend to remain fixated. Since Concept requires interpreting strategic linguistic signals within a shared conceptual space, success depends not only on linguistic fluency but also on adaptive exploration—an area where humans still hold a clear advantage.
Both for the static and dynamic prompt, we observe instances where the model repeats its previous (incorrect) answer, even though the prompt is updated with this information. Additionally, there are rounds where the guesses are not valid given the constraints of the highest level clue: for instance, if the first clue is ‘building’, the LLM’s guesses are not buildings.
Again contrary to our initial expectations, the reasoning model we included in our experiments (GPT-OSS-120B) does not outperform non-reasoning models. In some cases, this is caused because the number of reasoning tokens (n=1,000) were insufficient and the model did not arrive at an answer. This finding agrees with previous results that longer reasoning processes, in which the model generates more intermediate tokens, does not necessarily help the model to arrive at useful associations or correcting the erroneous directions (Yuan et al., 2025; Hassid et al., 2025; Vamvourellis and Mehta, 2025).
Table 3 shows the comparison of human and LLM performance per category (grouped by the first, highest-level clue). To compare LLMs with humans across categories, we extended the GLMM model to include solver (human vs. LLM), high-level-clue category, and their interaction as fixed effects. This allowed us to test whether differences in success rates between humans and LLMs varied systematically by concept category. For most models, there is no significant difference, except for Qwen2 72B (worse on ‘expression/quote, speak, word’, and ‘theater/film/camera’), LlaMA4-Maverick (worse on ‘expression/quote, speak, word’), and Mistral-Small 24B (worse on ‘expression/quote, speak, word’, and ‘object/thing/package’). To some extent, this agrees with our initial hypothesis that LLMs would struggle more with infrequent and culturally-dependent concepts, such as expressions and quotes.
Finally, we observe similar trends for Dutch, French, and Spanish, although the differences are even more pronounced. While human performance remains consistently high (94% found rate), all LLMs perform worse, and the gap relative to humans is larger than in English (see Table 4). Between the non-English languages, models seem to perform better on Spanish. This could be a reflection of common internet language distributions, where Spanish has a larger share compared to the other two languages. However, the models’ documentation does not provide detailed frequency distributions for non-English training data. Most model specifications list French and Spanish as supported languages, but not Dutch, yet LLMs’ performance on Dutch is not significantly worse than on French. This is particularly surprising for Mistral, given its French origins. Overall, the weak performance of LLMs on non-English Concept games highlights their limited capacity to transfer knowledge for this type of abstraction.
@@
Table 4: Found rates per model for French, Dutch, and Spanish. Since the static prompt consistently outperforms the dynamic one, we only include results for the static prompt. Entire dataset French: n=1,036, filtered: n=829. Entire dataset Dutch: n=1,155, filtered: n=932. Entire dataset Spanish: n=942, filtered: n=716.
5 Conclusion
In this paper, we present the game ‘Concept’ as a test for abductive reasoning. This game is comparable to Pictionary, but relying solely on natural language: Player 1 selects clues from a predefined set to describe a concept, Player 2 has to guess this concept. Whereas previous reasoning challenges such as ARC and SPaRC have addressed shortcomings in LLMs’ reasoning capacities, Concept levels the playing field by having the task representations rely only on natural language, which is better aligned with the pre-training data of LLMs.
We collected game logs in English, Dutch, French, and Spanish from the online boardgame platform Board Game Arena, providing a valuable human baseline performance. Per language, this results in about 1,000 concepts. We released the data for research purposes.
In our experiments, we prompted 7 state-of-the-art LLMs, both open- and closed-source. We implemented two prompting strategies, static and dynamic. The former presents the final set of clues given by Player 1, while the latter provides a level of interaction by presenting the clues iteratively, as they were given in the original game.
This language game, which is very easy for humans (with success rates over 90%), reveals important shortcomings of current LLMs. Specifically, we confirm previous findings that LLMs struggle to interpret Player 1’s strategic intents when giving the clues, which is supported by the finding that the dynamic prompt does not result in better results. Further, LLMs struggle to update their initial interpretations of the clues when sequential information updates are given. Even though incorrect guesses made by the model are added to the prompt, this does not help the model to arrive at the correct answer: often, the model remains stuck in the wrong direction. Additionally, we noted that in cases where both the LLM and humans identified the concept, LLMs consistently require fewer guesses, but in cases where more guesses are needed, humans are better at correcting their abstractions to still find the correct answer.
For further research, we suggest fine-tuning an LLM for this task, because previous research has shown that fine-tuning a small set of parameters can drastically increase performance on abstract reasoning tasks such as ACRE and RAVEN (Yun et al., 2025). Further, for interpretability, we suggest investigating the link between found rate of a concept and its frequency in the LLM’s pre-training data, or singling out the importance of one clue by implementing an ablation study. Additionally, since Concept can be played with multiple people guessing, it would be interesting to implement this as a multi-agent setting, in which multiple models guess, and each model sees the output of the others. Given the low correlation between models’ found rates, the interaction between different models could potentially aid performance.
Limitations
Acknowledgments
Appendix A Model specifications
In Table 5, we detail model name, version, and release date of the models included in our experiments.
Model Full Name / Variant Size Release Date Source
Appendix B Prompt
Below, we include the template for the static prompt in English:
[INST]Let’s play Concept! I will give you clues, you have to guess the concept. Only output the concept, nothing else! The hints are organized by color. Per color, the hierarchy of the clues is indicated by ’High-level’ and ’Low-level’. Low-level clues are given in function of the high-level clue of the same color. I will also give you the incorrect guesses, so you can learn from them. Here is an example: **Hints**: We are looking for a concept in this category: Building, Construction, City (highest-level hint, color: green). An additional hint in color green is: Metal. A high-level hint in color red is: Location, Country, Flag. An additional hint in color red is: Blue. An additional hint in color red is: White. An additional hint in color red is: Red. **Concept**: Eiffel Tower
Here are the clues: **Hints** [insert clues here] **Wrong guesses** [insert previous wrong answers here] **Concept**[/INST]
[INST]Let’s play Concept! I will give you clues, you have to guess the concept. Only output the concept, nothing else! The hints are organized by color. Per color, the hierarchy of the clues is indicated by ’High-level’ and ’Low-level’. Low-level clues are given in function of the high-level clue of the same color. I will also give you the incorrect guesses, so you can learn from them. Here is an example: **Hints**: We are looking for a concept in this category: Building, Construction, City (highest-level hint, color: green). An additional hint in color green is: Metal. A high-level hint in color red is: Location, Country, Flag. An additional hint in color red is: Blue. An additional hint in color red is: White. An additional hint in color red is: Red. **Concept**: Eiffel Tower
Here are the clues: **Hints** [insert clues until the current step here] **Wrong guesses** [insert previous wrong answers from LLMs and humans until the previous step here] **Concept**[/INST]
The static prompt in Spanish:
[INST]¡Juguemos a Concept! Te daré pistas en español y tú tendrás que adivinar el concepto en español. ¡Solo da el concepto, nada más! Las pistas están clasificadas por colores. Para cada color, la jerarquía de las pistas se indica con ’Nivel alto’ o ’adicional’. Las pistas adicionales se dan en función de las pistas de alto nivel del mismo color. También te daré las respuestas incorrectas para que puedas aprender de ellas. Aquí tienes un ejemplo: *Pistas**: Buscamos un concepto en esta categoría: Edificio, Construcción, Ciudad (pista de nivel superior, color: verde). Una pista adicional en color verde es: Metal. Una pista de alto nivel en color rojo es: Ubicación, País, Bandera. Una pista adicional en color rojo es: Azul. Una pista adicional en color rojo es: Blanco. Una pista adicional en color rojo es: Rojo. **Concepto**: Torre Eiffel.
Estas son las pistas: **Pistas** [insert clues here] **Respuestas incorrectas**: [insert previous wrong answers here] **Concepto**[/INST]
The static prompt in French:
[INST]Jouons à Concept! Je vais vous donner des indices en français, vous devez deviner le concept en français. Ne donnez que le concept, rien d’autre ! Les indices sont classés par couleur. Pour chaque couleur, la hiérarchie des indices est indiquée par « Haut niveau » et « Supplémentaire ». Les indices supplémentaires sont donnés en fonction de l’indice de haut niveau de la même couleur. Je vous donnerai également les réponses incorrectes, afin que vous puissiez en tirer des leçons. Voici un exemple: **Indices**: Nous recherchons un concept dans cette catégorie: Bâtiment, Construction, Ville (indice de plus haut niveau, couleur : vert). Un indice supplémentaire de couleur vert est: Métal. Un indice de haut niveau de couleur rouge est: Lieu, Pays, Drapeau. Un indice supplémentaire de couleur rouge est: Bleu. Un indice supplémentaire de couleur rouge est: Blanc. Un indice supplémentaire de couleur rouge est: Rouge. **Concept**: Tour Eiffel
Voici les indices: **Indices** [insert clues here] Réponses incorrectes**: [insert previous wrong answers here] **Concept**[/INST]
The static prompt in Dutch:
[INST]Laten we Concept spelen! Ik ga je aanwijzingen geven in het Nederlands, je moet het concept in het Nederlands raden. Geef alleen het concept, niets anders! De hints zijn gecategoriseerd op kleur. Voor elke kleur is de hiërarchie van de hints aangegeven door ’Hoog niveau’ of ’Extra’. De extra hints worden gegeven in functie van de hints op hoog-niveau van dezelfde kleur. Ik zal je ook de foute antwoorden geven, zodat je hiervan kunt leren. Hier is een voorbeeld: **Hints**: We zoeken een concept in deze categorie: Gebouw, Bouw, Stad (hoogste niveau hint, kleur: groen). Een extra hint in kleur groen is: Metaal. Een hint op hoog niveau in kleur rood is: Locatie, Land, Vlag. Een extra hint in kleur rood is: Blauw. Een extra hint in kleur rood is: Wit. Een extra hint in kleur rood is: Rood. **Concept**: Eiffeltoren
Hier zijn de hints: **Hints** [insert clues here] **Foute antwoorden**: [insert previous wrong answers here] **Concept**[/INST]

### v2

@@ 3 $answer\leftarrow\text{GENERATE\_ANSWER}(C+W)$;
However, as indicated by prior research, including the interaction between players gives more insights into Theory of Mind capabilities. Specifically, in Concept, Player 2 should keep in mind the strategic intent of Player 1’s choice of clues, and why Player 1 might choose to update their clues given Player 2’s incorrect answers. Therefore, in the dynamic prompt, we prompt iteratively. Specifically, we update the clues following the original game play. We present the clues per step (one step = the clues given until the original (human) Player 2 interrupted Player 1 by making a guess). For each step, we format the clues as mentioned earlier, and allow the model to guess as many times as humans did at that step. Then, instead of only adding the LLM’s incorrect answers, we also provide the human’s incorrect answers to approximate the original interactive process. Then, if the LLM has not found the correct answer, we move to the next step, updating the clues, and repeating the process. We formalize the procedure in pseudocode 2:
Input: $C_{1},\dots,C_{T}$ – clue sets for steps $t=1,\dots,T$
Input: $H_{1},\dots,H_{T}$ – human incorrect answers at each step (lists)
$W\leftarrow[\ ]$;
$\triangleright$ global list of incorrect answers
1 for $t\leftarrow 1$ to $T$ do
 $W.\text{extend}(H_{t})$;
 $\triangleright$ add human wrong answers at step $t$
2 for $k\leftarrow 1$ to $\lvert H_{t}\rvert$ do
3    $answer\leftarrow\text{GENERATE\_ANSWER}(C_{t}+W)$;
4    if IS_CORRECT(answer)(answer) then
5       return answeranswer;
6
    $W.\text{append}(answer)$;
    $\triangleright$ record model’s wrong answer
7
Algorithm 2 Dynamic prompt
@@ 3    $answer\leftarrow\text{GENERATE\_ANSWER}(C_{t}+W)$;
For the multilingual prompting, we translated the English prompt to the target language [^5] Each translation was proofread by at least a CEFR B2 (ACTFL Advanced Mid) speaker.. We compared translating the entire prompt to the target language, and keeping the instructions in English on a validation set (vatsal2025multilingualpromptengineeringlarge). Given that the latter yielded better results, on the test set we prompt with the entire prompt in the target language. We include all prompts in Appendix B.
@@ 3.3 Evaluation
We evaluate model performance using precision at k (prec@k), where a prediction is considered correct only if the guessed concept exactly matches the target concept. To ensure fair comparison, we apply extensive normalization to the LLM outputs, including correcting misrepresented characters (e.g., ASCII artifacts in non-English concepts), and cleaning responses such as removing the prefix ‘**Concept**:’ when present. For the static prompt setting, we fix $k$ at 10. For the dynamic prompt setting, $k$ is set to the number of guesses made by humans in the corresponding step (see Figure 2).
@@ 4 Results
Table 3: Found rates per model for French, Dutch, and Spanish for the static prompt. Entire dataset French: n=1,036, filtered: n=829. Entire dataset Dutch: n=1,155, filtered: n=932. Entire dataset Spanish: n=942, filtered: n=716.
4.1 Human performance
It is clear that Concept is an easy task for humans. To some extent, for persons who play the game very frequently, there might be a learning effect, but since there are 990 unique concepts per language in the game and our dataset includes 100 unique game plays, we can assume this effect is negligible. Considering all rounds, humans achieve a 92% found rate, and 93% on the filtered rounds. 63% of the rounds were solved within 10 guesses, although the correct answer was often still reached when more guesses were required. As expected, there was a negative correlation (Pearson’s r=0.3r=-0.3) between the number of clues provided and the likelihood of success, which could partially be explained because more difficult concepts requiring more clues. To examine whether performance varied across types of concepts, we categorized rounds by their first (and most important) clue (e.g., wildlife/animal, object/thing/package, building/construction/city, music/song/note), focusing on the top 20 most frequent categories. A Bayesian mixed-effects logistic regression model (GLMM) [^6] https://bambinos.github.io/bambi/notebooks/getting_started.html#generalized-linear-mixed-models was then fit with ‘found’ (i.e., whether the concept was found) as the outcome and category as a predictor [^7] Posterior estimates were obtained with NUTS sampling (4 chains; 1,000 warmup and 1,000 post-warmup draws), and category effects were considered meaningful when their 95% highest-density intervals excluded 0.. The analysis revealed no meaningful differences across categories, indicating that human performance was consistently high regardless of concept type. We refer to Table 5 in Appendix C for a complete overview of the results.
4.2 LLM performance
In comparison, all LLMs perform significantly worse than humans (see Table 2). Interestingly, in rounds where both humans and the LLM found the correct concept, the LLM overall needs fewer guesses to do so, suggesting greater efficiency once they are on the right track (see Figure 3).
Figure 3: Number of guesses needed to arrive at the correct concept, for concepts that are found both by humans and the LLM.
However, in rounds where the correct answer is not found easily, LLMs often get stuck pursuing wrong directions, whereas humans are better at revising their earlier interpretations and still arriving at the correct answer [^8] For models, we observe a strong negative correlation between found rate and number of guesses (Pearson’s $r$ between -0.68 and -0.74), whereas for humans the correlation is weak (r=0.14r=-0.14).. This finding echoes Chollet’s chollet2019measureintelligence view of exploration as a key component of intelligence: humans can flexibly revise their hypotheses, but LLMs tend to remain fixated. Since Concept requires interpreting strategic linguistic signals within a shared conceptual space, success depends not only on linguistic fluency but also on adaptive exploration—an area where humans still hold a clear advantage.
4.2.1 Correlation with human performance
Further, we compare human and LLM performance per category (grouped by the first, highest-level clue), by extending the GLMM model to include solver (human vs. LLM), high-level-clue category, and their interaction as fixed effects. This allowed us to test whether differences in success rates between humans and LLMs varied systematically by concept category. For most models, there is no significant difference, except for Qwen2 72B (worse on ‘expression/quote, speak, word’, and ‘theater/film/camera’), LlaMA4-Maverick (worse on ‘expression/quote, speak, word’), and Mistral-Small 24B (worse on ‘expression/quote, speak, word’, and ‘object/thing/package’). To some extent, this agrees with our initial hypothesis that LLMs would struggle more with infrequent and culturally-dependent concepts, such as expressions and quotes. We refer to Table 5 in Appendix C for a complete overview of the results.
4.2.2 Static vs. dynamic
A qualitative error analysis shows that both for the static and dynamic setting, models often repeat their previous (incorrect) answer, even though the prompt is updated with this information. Additionally, there are rounds where the guesses are not valid given the constraints of the highest level clue: for instance, if the first clue is ‘building’, the LLM’s guesses are not buildings.
4.2.3 Reasoning vs. non-reasoning models
Again contrary to our initial expectations, the reasoning model we included in our experiments (GPT-OSS-120B) does not outperform non-reasoning models. In some cases, this is caused because the number of reasoning tokens (n=1,000) were insufficient and the model did not arrive at an answer. This finding agrees with previous results that longer reasoning processes, in which the model generates more intermediate tokens, does not necessarily help the model to arrive at useful associations or correcting the erroneous directions (yuan2025turnaboutllmdeductivereasoningbenchmark; hassid2025dontoverthinkitpreferring; vamvourellis2025reasoningoverthinkingevaluatinglarge).
4.3 Multi-lingual performance
Finally, we observe similar trends for Dutch, French, and Spanish, although the differences are even more pronounced. While human performance remains consistently high (94% found rate), all LLMs perform worse, and the gap relative to humans is larger than in English (see Table 3). Between the non-English languages, models seem to perform better on Spanish. This could be a reflection of common internet language distributions, where Spanish has a larger share compared to the other two languages. Unfortunately, however, the models’ documentation does not provide detailed frequency distributions for non-English training data, so this remains a hypothesis. Most model specifications list French and Spanish as supported languages, but not Dutch, yet LLMs’ performance on Dutch is not significantly worse than on French. Overall, the weak performance of LLMs on non-English Concept games highlights their limited capacity to transfer knowledge for this type of abstraction.
5 Conclusion
In this paper, we present the word guessing game Concept as a test for abductive reasoning. Comparably to a single-modal version of Pictionary, Player 1 selects clues from a predefined set to describe a concept, Player 2 has to guess this concept. Whereas similar game-related benchmarks have addressed shortcomings in LLMs’ reasoning capacities, Concept levels the playing field by having the task representations rely only on natural language, which is better aligned with the pre-training data of LLMs.
We collected game logs in English, Dutch, French, and Spanish from the online boardgame platform Board Game Arena, providing a valuable human baseline performance.
To showcase the applicability of Concept, we provide baselines from 7 state-of-the-art LLMs, both open- and closed-source. This language game, which is very easy for humans (with success rates over 90%), reveals important shortcomings of current LLMs. Specifically, we confirm previous findings that LLMs struggle to interpret Player 1’s strategic intents when giving the clues, and fail to update their initial interpretations of the clues when sequential information updates are given. Even though incorrect guesses made by the model are added to the prompt, this does not help the model to arrive at the correct answer: often, the model remains stuck in the wrong direction. Additionally, we noted that in cases where both the LLM and humans identified the concept, LLMs consistently require fewer guesses, but in cases where more guesses are needed, humans are better at correcting their abstractions to still find the correct answer.
For further research, we suggest fine-tuning an LLM for this task. Further, for interpretability, we suggest investigating the link between found rate of a concept and its frequency in the LLM’s pre-training data, or singling out the importance of one clue by implementing an ablation study. Additionally, since Concept can be played with multiple people guessing, it would be interesting to implement this as a multi-agent setting, in which multiple models guess, and each model sees the output of the others. Given the low correlation between models’ found rates, the interaction between different models could potentially aid performance.
Limitations
Acknowledgments
Appendix A Model specifications
In Table 4, we detail model name, version, and release date of the models included in our experiments. For non-reasoning models, we allow 10 output tokens, and keep the temperature stable at 0.1 to keep the output more deterministic and reproducible. For GPT-OSS-120B, we allow 1,000 output tokens, with medium reasoning effort, and follow the recommended temperature setting (t=1). Because of hardware limitations, we prompt the models using the Together.ai API [^9] https://www.together.ai/, except for GPT 4.1 mini, for which we use the OpenAI API [^10] https://platform.openai.com/docs/models/gpt-4.1-mini. The total cost for prompting the models was 120\$.
Model Full Name / Variant Size Release Date Source
Appendix B Prompt
Below, we include the template for the static prompt in English:
[INST]Lets play Concept! I will give you clues, you have to guess the concept. Only output the concept, nothing else!
The hints are organized by color. Per color, the hierarchy of the clues is indicated by High-level and Low-level’. Low-level clues are given in function of the high-level clue of the same color. I will also give you the incorrect guesses, so you can learn from them. Here is an example: **Hints**:
We are looking for a concept in this category: Building, Construction, City (highest-level hint, color: green).
An additional hint in color green is: Metal.
A high-level hint in color red is: Location, Country, Flag.
An additional hint in color red is: Blue.
An additional hint in color red is: White.
An additional hint in color red is: Red.
**Concept**: Eiffel Tower
Here are the clues: **Hints** [insert clues here]
**Wrong guesses** [insert previous wrong answers here]
**Concept**[/INST]
[INST]Lets play Concept! I will give you clues, you have to guess the concept. Only output the concept, nothing else!
The hints are organized by color. Per color, the hierarchy of the clues is indicated by High-level and Low-level’. Low-level clues are given in function of the high-level clue of the same color. I will also give you the incorrect guesses, so you can learn from them. Here is an example: **Hints**:
We are looking for a concept in this category: Building, Construction, City (highest-level hint, color: green).
An additional hint in color green is: Metal.
A high-level hint in color red is: Location, Country, Flag.
An additional hint in color red is: Blue.
An additional hint in color red is: White.
An additional hint in color red is: Red.
**Concept**: Eiffel Tower
Here are the clues: **Hints** [insert clues until the current step here]
**Wrong guesses** [insert previous wrong answers from LLMs and humans until the previous step here]
**Concept**[/INST]
The static prompt in Spanish:
[INST]!‘Juguemos a Concept! Te daré pistas en espa~nol y tú tendrás que adivinar el concepto en espa~nol. !‘Solo da el concepto, nada más!
Las pistas están clasificadas por colores. Para cada color, la jerarquía de las pistas se indica con Nivel alto o adicional’. Las pistas adicionales se dan en función de las pistas de alto nivel del mismo color. También te daré las respuestas incorrectas para que puedas aprender de ellas.
Aquí tienes un ejemplo: **Pistas**:
Buscamos un concepto en esta categoría: Edificio, Construcción, Ciudad (pista de nivel superior, color: verde).
Una pista adicional en color verde es: Metal.
Una pista de alto nivel en color rojo es: Ubicación, País, Bandera.
Una pista adicional en color rojo es: Azul.
Una pista adicional en color rojo es: Blanco.
Una pista adicional en color rojo es: Rojo.
**Concepto**: Torre Eiffel.
Estas son las pistas: **Pistas** [insert clues here]
**Respuestas incorrectas**: [insert previous wrong answers here]
**Concepto**[/INST]
The static prompt in French:
[INST]Jouons à Concept! Je vais vous donner des indices en français, vous devez deviner le concept en français. Ne donnez que le concept, rien dautre !
Les indices sont classés par couleur. Pour chaque couleur, la hiérarchie des indices est indiquée par Haut niveau et Supplémentaire’. Les indices supplémentaires sont donnés en fonction de lindice de haut niveau de la m^eme couleur. Je vous donnerai également les réponses incorrectes, afin que vous puissiez en tirer des leçons.
Voici un exemple: **Indices**:
Nous recherchons un concept dans cette catégorie: B^atiment, Construction, Ville (indice de plus haut niveau, couleur : vert).
Un indice supplémentaire de couleur vert est: Métal.
Un indice de haut niveau de couleur rouge est: Lieu, Pays, Drapeau.
Un indice supplémentaire de couleur rouge est: Bleu.
Un indice supplémentaire de couleur rouge est: Blanc.
Un indice supplémentaire de couleur rouge est: Rouge.
**Concept**: Tour Eiffel
Voici les indices: **Indices** [insert clues here]
Réponses incorrectes**: [insert previous wrong answers here]
**Concept**[/INST]
The static prompt in Dutch:
[INST]Laten we Concept spelen! Ik ga je aanwijzingen geven in het Nederlands, je moet het concept in het Nederlands raden. Geef alleen het concept, niets anders!
De hints zijn gecategoriseerd op kleur. Voor elke kleur is de hiërarchie van de hints aangegeven door Hoog niveau of Extra’. De extra hints worden gegeven in functie van de hints op hoog-niveau van dezelfde kleur. Ik zal je ook de foute antwoorden geven, zodat je hiervan kunt leren.
Hier is een voorbeeld: **Hints**:
We zoeken een concept in deze categorie: Gebouw, Bouw, Stad (hoogste niveau hint, kleur: groen).
Een extra hint in kleur groen is: Metaal.
Een hint op hoog niveau in kleur rood is: Locatie, Land, Vlag.
Een extra hint in kleur rood is: Blauw.
Een extra hint in kleur rood is: Wit.
Een extra hint in kleur rood is: Rood.
**Concept**: Eiffeltoren
Hier zijn de hints: **Hints** [insert clues here]
**Foute antwoorden**: [insert previous wrong answers here]
**Concept**[/INST]
Appendix C Comparison per category
In Table 5 we give an overview of the results for each model per semantic category. The categories are defined as the highest-level (first) clue from the game, we analayze the 20 most frequent categories.